Diversification-Helen has 1/5 of her portfolio in U.S stocks, 1/8 of portfolio in European stocks, and 1/10 of the portfolio in Japanese stocks. The remainder is invested in muncipal bonds. What fraction of her portfolio is invested in the municipal bonds?
What percent is invested in muncipal bonds?

Answer by checkley77(12844)

1-(1/5+1/8+1/10)=1-(8+5+4)/40=1-17/40=23/40 was invested in municipal bonds.
23/40=.575=57.5% was invested in municipal bonds.